A “superfecta” is correctly guessing the top four horses in order in a horse race. (a)...

  1. A “superfecta” is correctly guessing the top four horses in order in a horse race.

    1. (a) If there are 8 horses in a race, how many different superfectas are possible?

    2. (b) “Boxing” a superfecta means that you get to choose the four horses for the superfecta without regard to order. How many box superfectas are possible?

    3. (c) If a bet on a superfecta costs 10 cents. How much would you expect for a box superfecta to cost? (Hint: Think about how many bets you are implicitly allowed to make by boxing!)
Answer #1

a) number of different superfectas(order matter) =(8P4) =8!/(8-4)! =8*7*6*5 =1680

b) number of box superfectas are possible =(8C4) =8!/(4!*(8-4)!)=70

c) expected cost for a box superfacta = 70*10 c = 700 cents or $7